A parallelogram has sides 4 cm. and 18 cm. and included angle 30º. Find one side

of the square of equal area.



1
Expert's answer
2021-02-02T05:08:05-0500

Area of the parallelogram: "A=4*18*sin30^0=36 \;cm^2."

Let the side of the square be x.

So, "x^2=36."

"x=\sqrt{36}=6 \;cm."